# Twisted Superfields

### v7

@@ 2 Dynamical Symmetry Breaking
$T^{Y}$ $={\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}{\rm diag}\left({\tfrac{2}{3}},{\tfrac{2}{3}},{\tfrac{2}{3}},1,1,0\right)$ (2.15)
$T_{1}^{Y}$ $=-{\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}{\rm diag}\left({\tfrac{4}{3}},{\tfrac{4}{3}},{\tfrac{4}{3}},1,1,2\right)$
$T_{3}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(-1,-1,-1,1,1,0\right)$
$T_{1}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(2,2,2,-1,-1,-2\right)$
$T^{Y}$ is the $A_{\mu}^{Y}$ group structure when acting on gaugino, adjoint, flavor 2 or flavor 3 fields. $T_{1}^{Y}$ is the $A_{\mu}^{Y}$ group structure when acting on flavor 1 fields. $T_{3}^{Y^{\prime}}$ and $T_{1}^{Y^{\prime}}$ are the $A_{\mu}^{Y^{\prime}}$ group structures when acting on flavor 3 and 1 fields. The $A_{\mu}^{Y^{\prime}}$ group structures when acting on gauginos, adjoint, or flavor 2 chiral fields can be calculated from eqs (1.20), (2.13) and (2.14).
It will be seen below that the $T^{Y}$ and $T_{1}^{Y}$ matrices have the correct form for their gauge boson $A_{\mu}^{Y}$ to be identified as the U(1) weak hypercharge field of the Standard Model with $-{\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}g_{Y}$ identified as the weak hypercharge coupling. The $A_{\mu}^{Y^{\prime}}$ gauge field with its coupling $g_{Y^{\prime}}$ is a second U(1) gauge boson in this model that remains massless at the unification scale. Due to eq (2.11), there is no mixing between the $A_{\mu}^{Y}$ and $A_{\mu}^{Y^{\prime}}$ gauge fields. Below the unification scale, the couplings for the SU(3), SU(2), $Y$ and $Y^{\prime}$ groups run differently, so they are denoted by $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$.
@@
$\left(\begin{array}[]{c}{eT_{F}^{\gamma}}\\ {g_{Z^{\prime}}T_{F}^{Z^{\prime}}}\\ {g_{Z}T_{F}^{Z}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{Z}}&{\sin\phi_{Z}}\\ {0}&{-\sin\phi_{Z}}&{\cos\phi_{Z}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{Z}}&{0}&{\sin\theta_{Z}}\\ {0}&{1}&{0}\\ {-\sin\theta_{Z}}&{0}&{\cos\theta_{Z}}\end{array}\right)\left(\begin{array}[]{c}{g_{Y}T_{F}^{Y}}\\ {g_{Y^{\prime}}T_{F}^{Y^{\prime}}}\\ {g_{2}T_{2}^{3}}\end{array}\right).$ (2.16)
The angle $\theta_{Z}$ is chosen to make $(T_{1}^{\gamma})_{44}=(T^{\gamma})_{55}=0$, so that the photon gets no mass from $\bar{\phi}_{21}$, $\bar{\tilde{\phi}}_{22}$ or $\bar{\varphi}_{2}$.
The angle $\phi_{Z}$ is chosen to make $(T_{1}^{Z^{\prime}})_{44}=0$ and $(T_{A}^{Z^{\prime}})_{55}=(T_{A}^{Z^{\prime}})_{66}$, so that the $Z^{\prime}$ gets no mass from $\bar{\phi}_{21}$ or $\bar{\varphi}_{2}$. The resulting angles are given by:
@@
In section 4, it will be argued that $g_{Y^{\prime}}$ will be driven to a very small value from an effective anomaly generated by a very massive right-handed neutrino. As a result, the angle $\phi_{Z}$ is very small, and $\theta_{Z}$ is very close to the Weinberg angle $\theta_{W}$ of the Standard Model.
With the following coupling constant normalization,
$e=-g_{2}\sin\theta_{Z}\,,$ (2.18)
@@
$eT^{\gamma}$ $=-e{\rm diag}\left({\tfrac{1}{3}},{\tfrac{1}{3}},{\tfrac{1}{3}},1,0,0\right)$ (2.19)
$eT_{1}^{\gamma}$ $=e{\rm diag}\left({\tfrac{2}{3}},{\tfrac{2}{3}},{\tfrac{2}{3}},0,1,1\right)\,.$
@@
$g_{Z}T^{Z}$ $\simeq-(g_{2}/\cos\theta_{W}){\rm diag}\left({\tfrac{1}{3}x},{\tfrac{1}{3}x},{\tfrac{1}{3}x},-\tfrac{1}{2}+x,\tfrac{1}{2},0\right)$ (2.20)
$g_{Z}T_{1}^{Z}$ $\simeq(g_{2}/\cos\theta_{W}){\rm diag}\left({\tfrac{2}{3}x},{\tfrac{2}{3}x},{\tfrac{2}{3}x},\tfrac{1}{2},-\tfrac{1}{2}+x,x\right)$
$|\phi_{Z}|$ $\ll 1,$
where $x=\sin^{2}\theta_{Z}\simeq\sin^{2}\theta_{W}$. The group structure of the $Z^{\prime}$ boson is shown in eq (5.1) of section 5.
@@
$\psi_{11}=\left(\begin{array}[]{c}{u_{1}^{\left(1\right)}}\\ {u_{2}^{\left(1\right)}}\\ {u_{3}^{\left(1\right)}}\end{array}\right)\,\psi_{12}=\left(\begin{array}[]{c}{d_{1}^{\left(2\right)}}\\ {d_{2}^{\left(2\right)}}\\ {d_{3}^{\left(2\right)}}\end{array}\right)\,\psi_{13}=\left(\begin{array}[]{c}{d_{1}^{\left(3\right)}}\\ {d_{2}^{\left(3\right)}}\\ {d_{3}^{\left(3\right)}}\end{array}\right)\,\psi_{21}=\left(\begin{array}[]{c}{\nu_{W}^{\left(1\right)}}\\ {e_{W}^{\left(1\right)}}\\ {e^{\left(1\right)}}\end{array}\right)\,\psi_{22}=\left(\begin{array}[]{c}{e_{W}^{\left(2\right)}}\\ {\nu_{W}^{\left(2\right)}}\\ {\nu^{\left(2\right)}}\end{array}\right)\,\psi_{23}=\left(\begin{array}[]{c}{e_{W}^{\left(3\right)}}\\ {\nu_{W}^{\left(3\right)}}\\ {\nu^{\left(3\right)}}\end{array}\right)$
$\tilde{\psi}_{11}^{T}=\left(\begin{array}[]{c}{\tilde{u}_{1}^{\left(1\right)}}\\ {\tilde{u}_{2}^{\left(1\right)}}\\ {\tilde{u}_{3}^{\left(1\right)}}\end{array}\right)\,\tilde{\psi}_{12}^{T}=\left(\begin{array}[]{c}{\tilde{d}_{1}^{\left(2\right)}}\\ {\tilde{d}_{2}^{\left(2\right)}}\\ {\tilde{d}_{3}^{\left(2\right)}}\end{array}\right)\,\tilde{\psi}_{13}^{T}=\left(\begin{array}[]{c}{\tilde{d}_{1}^{\left(3\right)}}\\ {\tilde{d}_{2}^{\left(3\right)}}\\ {\tilde{d}_{3}^{\left(3\right)}}\end{array}\right)\,\tilde{\psi}_{21}^{T}=\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{\left(1\right)}}\\ {\tilde{e}_{W}^{\left(1\right)}}\\ {\tilde{e}^{\left(1\right)}}\end{array}\right)\,\tilde{\psi}_{22}^{T}=\left(\begin{array}[]{c}{\tilde{e}_{W}^{\left(2\right)}}\\ {\tilde{\nu}_{W}^{\left(2\right)}}\\ {\tilde{\nu}^{\left(2\right)}}\end{array}\right)\,\tilde{\psi}_{23}^{T}=\left(\begin{array}[]{c}{\tilde{e}_{W}^{\left(3\right)}}\\ {\tilde{\nu}_{W}^{\left(3\right)}}\\ {\tilde{\nu}^{\left(3\right)}}\end{array}\right).$
@@
In the above labelling, lower numerical indices are fundamental-representation indices for the unbroken SU(3) group (the strong interaction). Fermions with a “W” index interact with the W boson (as members of an isodoublet). Based on the magnitude of their electric charges, up-type quarks, down-type quarks, charged leptons, and neutral leptons are labelled with $u,d,e,\nu$. All of the fermion fields are 2-component Weyl fermions with a lower, undotted spin index. In the convention of [7] (which is also the Wess/Bagger & Bilal convention of [15]), a Weyl fermion with a lower undotted index corresponds to a right-chiral fermion that vanishes when acted on by $1-\gamma_{5}$ (see appendix A of [16]).
@@
In that convention, the $u$ and $d$ fields are right-chiral fermions with electric charges of ${\tfrac{2}{3}}$ and $-{\tfrac{1}{3}}$, respectively, so they are mapped to right-chiral quarks. The fields $\tilde{u}$ and $\tilde{d}$ are right-chiral fermions with electric charges of $-{\tfrac{2}{3}}$ and ${\tfrac{1}{3}}$, respectively, so they are mapped to Hermitian conjugates of left-chiral quarks. One way that this model differs from the Standard Model is that some of the right-chiral quarks have a “W” index so they interact with the W boson, while some of the left-chiral quarks lack that index so they do not interact with the W boson. That difference from the Standard Model will be discussed further in the next section of this paper.
Mapping of leptons in this model is described in the next section, but one case is mentioned here: The right-chiral fields $e_{W}^{\left(1\right)}$ and $\nu_{W}^{\left(1\right)}$ (with electric charges of $+1$ and 00) are mapped to $e_{L}^{\dagger}$ and $\nu_{e}^{\dagger}$, Hermitian conjugates of the left-chiral electron and its neutrino. The right-chiral field $\tilde{e}^{\left(1\right)}$ (with charge $-1$ and no $W$ subscript) is mapped to $e_{R}$, the right-chiral electron. Since only fields with a $W$ subscript connect to the W boson, this convention ensures that only the left-chiral electrons (and their Hermitian conjugates) interact with the W boson.
@@ 3.1 Confinement
For a supersymmetric gauge theory involving a chiral superfield in the adjoint representation, it has been shown that a tree-level mass term $m_{\Phi}$ for the adjoint superfield will lead to quark confinement [17, 18, 15, 14]. The duality inherent in these theories permits moving from a description in terms of strongly coupled scalars with color-electric charge to a description in terms of weakly coupled monopoles with color-magnetic charge. A tree-level $m_{\Phi}$ mass term can cause the vacuum to settle on one of two configurations where the vev of the trace of the square of the adjoint superfield does not vanish $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle\neq 0$. In one of those configurations, color-magnetic monopoles become massless, condense, and cause quark confinement through a dual Meisner effect [17, 18, 15, 14].
The theory of this paper is not supersymmetric, since its “superfields” are “twisted”. Nonetheless, this theory does have an adjoint-representation “twisted superfield” with a mass $m_{\Phi}$. Also, the scalars in this theory are in the same representation as the scalars in the corresponding supersymmetric theory, so the vacuum moduli spaces of the two theories should be similar, particularly below the scale where the $m=1$ fundamental and conjugate scalars of the theory get large masses (see section 5). That being the case, it is speculated that quantum effects similar to those in the supersymmetric theory cause the following vev to form:
@@
It is further speculated that the similarity with the corresponding supersymmetric moduli space is sufficient so that the scalars in $\varphi_{1}$ form color-magnetic monopoles that become massless, condense and cause quark confinement through the dual Meisner effect.
@@
$2\sqrt{2}g_{3}i\int d^{4}x\rm{tr}\left(\tilde{\chi}\left\langle\varphi_{1}^{\dagger}\right\rangle\lambda-\tilde{\lambda}\left\langle\varphi_{1}^{\dagger}\right\rangle\chi\right)+h.c.,$ (3.2)
@@
Before concluding this subsection, there are a couple of additional quantum effects to be discussed. For supersymmetric theories, instantons can generate nonperturbative low-energy effective superpotential terms. The appendix proposes that a similar effect occurs for this theory. To accommodate these terms, it is assumed that the following vevs that are zero classically acquire small but nonzero vevs quantum mechanically:
$\bar{\tilde{\phi}}_{21},\bar{\phi}_{22}\neq 0\textrm{ quantum mechanically}.$ (3.3)
These vevs set the scale for neutrino masses later in this section.
As discussed in the appendix, the quantum-generated superpotential terms also try to make $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$ small. To accomplish this in the presence of eq (3.1), it is assumed that the quantum vacuum modifies $\left\langle\varphi_{2}\right\rangle$ of eq (2.6) to the following:
$\left\langle\varphi_{2}\right\rangle=\frac{i\bar{\varphi}_{2}}{\sqrt{2\left(1+c^{2}\right)}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{c}\\ {0}&{1}&{0}\end{array}\right),$ (3.4)
where the quantum-generated $c$ is assumed to be much smaller than 1. The square of the above vev has a nonvanishing trace that can reduce $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$.

### v8

@@ 2 Dynamical Symmetry Breaking
$T_{23}^{Y}=T_{22}^{Y}$ $={\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}{\rm diag}\left({\tfrac{2}{3}},{\tfrac{2}{3}},{\tfrac{2}{3}},1,1,0\right)$ (2.14)
$T_{21}^{Y}=T_{1F}^{Y}$ $=-{\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}{\rm diag}\left({\tfrac{4}{3}},{\tfrac{4}{3}},{\tfrac{4}{3}},1,1,2\right)$
$T_{23}^{Y^{\prime}}=T_{22}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(-1,-1,-1,1,1,0\right)$ (2.15)
$T_{21}^{Y^{\prime}}=T_{1F}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(2,2,2,-1,-1,-2\right)$
Either version of $T_{mF}^{Y}$ can be used when acting on gaugino or adjoint fields since $T_{+}^{0}$ commutes with those. The $A_{\mu}^{Y^{\prime}}$ group structures when acting on gauginos and adjoint fermions is the same as $T_{23}^{Y^{\prime}}$.
It will be seen below that the $T_{mF}^{Y}$ matrices have the correct form for their gauge boson $A_{\mu}^{Y}$ to be identified as the U(1) weak hypercharge field of the Standard Model with $-{\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}g_{Y}$ identified as the weak hypercharge coupling. The $A_{\mu}^{Y^{\prime}}$ gauge field with its coupling $g_{Y^{\prime}}$ is a second U(1) gauge boson in this model that remains massless at the unification scale. Due to eq (2.10), there is no mixing between the $A_{\mu}^{Y}$ and $A_{\mu}^{Y^{\prime}}$ gauge fields. Below the unification scale, the couplings for the SU(3), SU(2), $Y$ and $Y^{\prime}$ groups run differently, so they are denoted by $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$.
@@
$\left(\begin{array}[]{c}{eT_{mF}^{\gamma}}\\ {g_{Z^{\prime}}T_{mF}^{Z^{\prime}}}\\ {g_{Z}T_{mF}^{Z}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{Z}}&{\sin\phi_{Z}}\\ {0}&{-\sin\phi_{Z}}&{\cos\phi_{Z}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{Z}}&{0}&{\sin\theta_{Z}}\\ {0}&{1}&{0}\\ {-\sin\theta_{Z}}&{0}&{\cos\theta_{Z}}\end{array}\right)\left(\begin{array}[]{c}{g_{Y}T_{mF}^{Y}}\\ {g_{Y^{\prime}}T_{mF}^{Y^{\prime}}}\\ {g_{2}T_{2}^{3}}\end{array}\right).$ (2.16)
The angle $\theta_{Z}$ is chosen to make $(T_{21}^{\gamma})_{44}=(T_{22}^{\gamma})_{55}=0$, so that the photon gets no mass from $\bar{\phi}_{21}$, $\bar{\phi}_{22}$ or $\bar{\varphi}_{2}$.
The angle $\phi_{Z}$ is chosen to make $(T_{21}^{Z^{\prime}})_{44}=0$. The resulting angles are given by:
@@
In section 4, it will be argued that $g_{Z^{\prime}}$ will be driven to a very small value from an effective anomaly. As a result, the angle $\phi_{Z}$ is very small, and $\theta_{Z}$ is very close to the Weinberg angle $\theta_{W}$ of the Standard Model.
@@
$eT_{22}^{\gamma}=eT_{23}^{\gamma}$ $=-e{\rm diag}\left({\tfrac{1}{3}},{\tfrac{1}{3}},{\tfrac{1}{3}},1,0,0\right)$ (2.18)
$eT_{21}^{\gamma}=eT_{1F}^{\gamma}$ $=e{\rm diag}\left({\tfrac{2}{3}},{\tfrac{2}{3}},{\tfrac{2}{3}},0,1,1\right)\,,$
where the following coupling constant normalization is used:
$e=-g_{2}\sin\theta_{Z}\,.$ (2.19)
@@
$g_{Z}T_{22}^{Z}$ $\simeq-(g_{2}/\cos\theta_{W}){\rm diag}\left({\tfrac{1}{3}x},{\tfrac{1}{3}x},{\tfrac{1}{3}x},-\tfrac{1}{2}+x,\tfrac{1}{2},0\right)$ (2.20)
$g_{Z}T_{1F}^{Z}$ $\simeq(g_{2}/\cos\theta_{W}){\rm diag}\left({\tfrac{2}{3}x},{\tfrac{2}{3}x},{\tfrac{2}{3}x},\tfrac{1}{2},-\tfrac{1}{2}+x,x\right)$
$x$ $=\sin^{2}\theta_{Z}\simeq\sin^{2}\theta_{W}\,.$
The group structure of the $Z^{\prime}$ boson is shown in eq (5.2) of section 5.
@@
$\psi_{11}=\left(\begin{array}[]{c}{u_{1}^{\left(1\right)}}\\ {u_{2}^{\left(1\right)}}\\ {u_{3}^{\left(1\right)}}\end{array}\right)\,\psi_{12}=\left(\begin{array}[]{c}{d_{1}^{\left(2\right)}}\\ {d_{2}^{\left(2\right)}}\\ {d_{3}^{\left(2\right)}}\end{array}\right)\,\psi_{13}=\left(\begin{array}[]{c}{d_{1}^{\left(3\right)}}\\ {d_{2}^{\left(3\right)}}\\ {d_{3}^{\left(3\right)}}\end{array}\right)\,\psi_{21}=\left(\begin{array}[]{c}{\nu_{W}^{\left(1\right)}}\\ {e_{W}^{\left(1\right)}}\\ {e^{\left(1\right)}}\end{array}\right)\,\psi_{22}=\left(\begin{array}[]{c}{\nu_{W}^{\left(2\right)}}\\ {e_{W}^{\left(2\right)}}\\ {e^{\left(2\right)}}\end{array}\right)\,\psi_{23}=\left(\begin{array}[]{c}{\nu_{W}^{\left(3\right)}}\\ {e_{W}^{\left(3\right)}}\\ {e^{\left(3\right)}}\end{array}\right)$
$\tilde{\psi}_{11}^{T}=\left(\begin{array}[]{c}{\tilde{u}_{1}^{\left(1\right)}}\\ {\tilde{u}_{2}^{\left(1\right)}}\\ {\tilde{u}_{3}^{\left(1\right)}}\end{array}\right)\,\tilde{\psi}_{12}^{T}=\left(\begin{array}[]{c}{\tilde{d}_{1}^{\left(2\right)}}\\ {\tilde{d}_{2}^{\left(2\right)}}\\ {\tilde{d}_{3}^{\left(2\right)}}\end{array}\right)\,\tilde{\psi}_{13}^{T}=\left(\begin{array}[]{c}{\tilde{d}_{1}^{\left(3\right)}}\\ {\tilde{d}_{2}^{\left(3\right)}}\\ {\tilde{d}_{3}^{\left(3\right)}}\end{array}\right)\,\tilde{\psi}_{21}^{T}=\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{\left(1\right)}}\\ {\tilde{e}_{W}^{\left(1\right)}}\\ {\tilde{e}^{\left(1\right)}}\end{array}\right)\,\tilde{\psi}_{22}^{T}=\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{\left(2\right)}}\\ {\tilde{e}_{W}^{\left(2\right)}}\\ {\tilde{e}^{\left(2\right)}}\end{array}\right)\,\tilde{\psi}_{23}^{T}=\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{\left(3\right)}}\\ {\tilde{e}_{W}^{\left(3\right)}}\\ {\tilde{e}^{\left(3\right)}}\end{array}\right).$
@@
In the above labelling, lower numerical indices are fundamental-representation indices for the unbroken SU(3) group (the strong interaction). Fermions with a “W” index interact with the $W$ boson (as members of an isodoublet). Based on the magnitude of their electric charges, up-type quarks, down-type quarks, charged leptons, and neutrinos are labelled with $u,d,e,\nu$. All of the fermion fields are 2-component Weyl fermions with a lower, undotted spin index. In the convention of [6] (which is also the Wess/Bagger & Bilal convention of [14]), a Weyl fermion with a lower undotted index corresponds to a right-chiral fermion that vanishes when acted on by $1-\gamma_{5}$ (see appendix A of [15]).
@@
In that convention, the $u$ and $d$ fields are right-chiral fermions with electric charges of ${\tfrac{2}{3}}$ and $-{\tfrac{1}{3}}$, respectively, so they are mapped to right-chiral quarks. The fields $\tilde{u}$ and $\tilde{d}$ are right-chiral fermions with electric charges of $-{\tfrac{2}{3}}$ and ${\tfrac{1}{3}}$, respectively, so they are mapped to Hermitian conjugates of left-chiral quarks. One way that this model differs from the Standard Model is that some of the right-chiral quarks have a “$W$” index so they interact with the $W$ boson, while some of the left-chiral quarks lack that index so they do not interact with the $W$ boson. That difference is discussed in the next section.
@@ 3.1 Confinement
For a supersymmetric gauge theory involving a chiral superfield in the adjoint representation, it has been shown that a tree-level mass term $m_{\Phi}$ for the adjoint superfield will lead to quark confinement [16, 17, 14, 13]. The duality inherent in these theories permits moving from a description in terms of strongly coupled scalars with color-electric charge to a description in terms of weakly coupled monopoles with color-magnetic charge. A tree-level $m_{\Phi}$ mass term can cause the vacuum to settle on one of two configurations where the vev of the trace of the square of the adjoint superfield does not vanish $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle\neq 0$. In one of those configurations, color-magnetic monopoles become massless, condense, and cause quark confinement through a dual Meissner effect [16, 17, 14, 13].
The theory of this paper is not supersymmetric, since its “superfields” are “twisted”. Nonetheless, this theory does have an adjoint-representation “twisted superfield” with a mass $m_{\Phi}$. Also, the scalars in this theory are in the same representation as the scalars in the corresponding supersymmetric theory, so the vacuum moduli spaces of the two theories should be similar, particularly below the scale where the $m=1$ fundamental and conjugate scalars of the theory (aka leptoquarks) get large masses (see section 5). That being the case, it is speculated that quantum effects similar to those in the supersymmetric theory cause the following vev to form:
@@
It is further speculated that the similarity with the corresponding supersymmetric moduli space is sufficient so that the scalars in $\varphi_{1}$ form color-magnetic monopoles that become massless, condense and cause quark confinement through the dual Meissner effect.
@@
$-\sqrt{2}g_{3}i\int d^{4}x\rm{tr}\left(\tilde{\chi}\left\langle\varphi_{1}^{\dagger}\right\rangle\lambda-\tilde{\lambda}\left\langle\varphi_{1}^{\dagger}\right\rangle\chi\right)+h.c.,$ (3.2)
@@
There are a couple of additional quantum effects pertinent to this section. For supersymmetric theories, instantons can generate nonperturbative low-energy effective superpotential terms. The Appendix proposes that a similar effect occurs for this theory. To accommodate these terms, the following vevs that are zero classically acquire small but nonzero vevs quantum mechanically:
$\bar{\tilde{\phi}}_{21},\bar{\tilde{\phi}}_{22}\neq 0\textrm{ quantum mechanically}.$ (3.3)
A quantum superpotential term proportional to $1/\bar{\tilde{\phi}}_{21}$ (see eq (A.13)) sets the scale for masses of the heavy right-handed neutrinos discussed later in this section.
As discussed in the Appendix, the quantum-generated superpotential terms also try to reduce the value of $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$. To accomplish this in the presence of eq (3.1), the quantum vacuum modifies $\left\langle\varphi_{2}\right\rangle$ of eq (2.6) to the following:
$\left\langle\varphi_{2}\right\rangle=\frac{\bar{\varphi}_{2}}{\sqrt{2\left(1+c^{2}\right)}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{1}\\ {0}&{c}&{0}\end{array}\right),$ (3.4)
where the quantum-generated $c$ is much smaller than 1. The square of the above vev has a nonvanishing trace that can reduce $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$.